OpenStudy (anonymous):

y=1/3(x-1/2)-1/3 In Standard Form Ax+By=C

OpenStudy (anonymous):

y=1/3x-1/6-1/3 2x-6y=3

OpenStudy (anonymous):

I know that

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

so, what do you mean

OpenStudy (anonymous):

i'm waiting to you

OpenStudy (anonymous):

ok how do you get there from y=1/3x -2

OpenStudy (anonymous):

no but y=1/3x-1/2

OpenStudy (anonymous):

y = 3/(3x - 3/2) - 1/3 [distributive law applied to denominator] y(3x - 3/2) = 3 - (1/3)(3x - 3/2) [multiplying both sides by (3x - 3/2) ] 6y(3x - 3/2) = 6(3 - (1/3)(3x - 3/2) [multiplying both sides by 6 to clear of fractions] y(18x - 9) = 18 - 2(3x - 3/2) [distributive law] y(18x - 9) = 18 -6x + 3 [distributive law] y(18x - 9) = -6x + 21 [combining like terms] y = (-6x + 21)/(18x - 9)[dividing both sides by (18x - 9) ] y = (-2x + 7)/(6x - 3). This is the equation of a hyperbola, not a straight line. Consequently, no such form exists for the given equation.